OpenStudy (anonymous):

Select the factors of ax − 2ay + 3bx − 6by.

OpenStudy (callisto):

ax − 2ay + 3bx − 6by =a(x-2y) +3b (x-2y) =(a+3b)(x-2y)

OpenStudy (hoblos):

ax − 2ay + 3bx − 6by = a(x-2y) +3b(x-2y) = (x-2y)(a+3b)

OpenStudy (tommo_lcfc):

This expression can be simplified to give \[a(x-2y) + 3b(x-2y)\] which is \[(a+3b)(x-2y)\]